OpenStudy (cutiekawdfi):

identify the conic section, lines of symmetry, and the domain and range. 1. x^2+y2=4 2. x=2y^2+5

OpenStudy (compassionate):

First make the equations equal 0.

OpenStudy (zarkon):

\[x^2+y^2=4\] is the same as \[(x-0)^2+(y-0)^2=2^2\]